a 10.00 gram sample of the hydrated salt cacl2. x.h2o was thermally decomposed by anhydrous calcium chloride and water vapours find the no. of water molecules in the salt


1
Expert's answer
2021-05-05T07:09:16-0400

moles of Cacl2 "= \frac{ 7.93}{110.98} = 0.0715 mole."



The loss of water amounts to 10 g - 7.93 g = 2.07 grams of water.


H2O moles "=\frac{2.07 }{ 18 } = 0.115 mole."


The ratio of water to cacl2 (which is what ‘x’ means in K2CO3xH2O) is equal to:


Ratio "=\frac {0.115}{0.0715}=1.6 = 2"


So, x is 2 and the formula is "CaCl_2. 2H_2O"